# Discovering novel order parameters in the Potts model: A bridge between machine learning and critical phenomena

### v1

@@ III.1 $q=3$ Potts model on a square lattice
Table 1: Critical properties of the 3-state Potts model at $d=2$ on a square lattice, obtained using the alternative order parameters $M_{\mathrm{max}}$, $M_{\mathrm{med}}$, $M_{\mathrm{min}}$ and $M_{\mathrm{map}}$. The critical temperature $T_{c}$, scaling dimension $\Delta_{\sigma}$, and critical exponent $\nu$, along with their standard errors, are estimated via finite-size scaling within the interval $T\in(0.920,1.046)$ with step size $\Delta T=0.002$. Theoretical values from the literature are included for comparison.
@@ III.2 $q=4$ Potts model on a square lattice
For the $q=4$ case, besides the most and least states, we also have the second most and second least states for each configuration, therefore we define
@@
and $M_{\mathrm{map}}$ is defined the same as in Eq. (7). As the analysis for $3$-state Potts model, we assume that in one configuration, the occupancy of four states descends as $a,b,c,d$. With the Z(4) symmetry, we could have 24 cases of how these four occupancies are assigned to these $4$ states. After smearing two specific states of these four ones with the mapping procedure, the average $M_{\mathrm{map}}$ of these 24 configurations equals $(3(a-d)+(b-c))/6$. Due to that $M_{\mathrm{max}}=(4a-1)/3$ and $M_{\mathrm{min}}=(4d-1)/3$ as in Eqs. (3) and (8) and $M_{\mathrm{max}}^{\mathrm{sec}}=(4b-1)/3$, and $M_{\mathrm{min}}^{\mathrm{sec}}=(4c-1)/3$ as in Eqs. (11) and (12), we derive the fundamental relationship
@@
where the line signifies the ensemble average. This is also confirmed numerically in Fig. 3, where $M_{\mathrm{max}}$, $M_{\mathrm{max}}^{\mathrm{sec}}$, $M_{\mathrm{min}}^{\mathrm{sec}}$, $M_{\mathrm{min}}$ and $M_{\mathrm{map}}$ as functions of temperature $T$ with lattice size $L=80$ are shown and the curve of $M_{\mathrm{map}}$ labeled with red triangles and that of $\frac{1}{8}(3(M_{\mathrm{max}}-M_{\mathrm{min}})+M_{\mathrm{max}}^{\mathrm{sec}}-M_{\mathrm{min}}^{\mathrm{sec}})$ labeled with gray dots collapse on top of each other. Again, it’s worth noting that the fluctuations of $M_{\mathrm{map}}$ are much larger than those of $\frac{1}{8}(3(M_{\mathrm{max}}-M_{\mathrm{min}})+M_{\mathrm{max}}^{\mathrm{sec}}-M_{\mathrm{min}}^{\mathrm{sec}})$, which is mainly due to the fluctuations between those 24 configurations for $M_{\mathrm{map}}$.
Figure 3: (Color online) Alternative order parameters $M_{\mathrm{max}}$, $M_{\mathrm{max}}^{\mathrm{sec}}$, $M_{\mathrm{min}}^{\mathrm{sec}}$, $M_{\mathrm{min}}$ and $M_{\mathrm{map}}$ as functions of temperature $T$ for the $4$-state Potts model at $d=2$, simulated on a square lattice of size $L=80$. The relation $M_{\mathrm{map}}=\frac{1}{8}(3(M_{\mathrm{max}}-M_{\mathrm{min}})+M_{\mathrm{max}}^{\mathrm{sec}}-M_{\mathrm{min}}^{\mathrm{sec}})$ is confirmed by the overlap of the curves. The blue horizontal dashed line marks zero to indicate the sign, while the red vertical dashed line denotes the theoretical critical temperature $T_{c}$.
Order parameter $T_{c}$ $\Delta_{\sigma}$ $\nu$
@@
Theoretical value 1/$\ln{(1+\sqrt{4})}\sim 0.9102$ 0.128* 0.722*
Table 2: Critical properties of the 4-state Potts model at $d=2$ on a square lattice, obtained using the alternative order parameters $M_{\mathrm{max}}$, $M_{\mathrm{max}}^{\mathrm{sec}}$, $M_{\mathrm{min}}^{\mathrm{sec}}$, $M_{\mathrm{min}}$ and $M_{\mathrm{map}}$. The critical temperature $T_{c}$, scaling dimension $\Delta_{\sigma}$, and critical exponent $\nu$, along with their standard errors, are estimated via finite-size scaling within the interval $T\in(0.840,0.966)$ with step size $\Delta T=0.002$. Theoretical values from the literature are included for comparison and the values of $\Delta_{\sigma}$ and $\nu$ marked with * are without higher-order corrections.
The temperature dependence of $M_{\mathrm{max}}$, $-M_{\mathrm{max}}^{\mathrm{sec}}$, $|M_{\mathrm{min}}^{\mathrm{sec}}|$, $|M_{\mathrm{min}}|$ and $M_{\mathrm{map}}$ and their associated finite-size scaled results are shown for better visualization, respectively, in Figs. A4-A8 in Appendix with the known Monte-Carlo numerical value of $\nu=0.722$ and $\Delta_{\sigma}=0.128$ [56], which validates that all of these observables could serve as alternative order parameters. With the FSS procedure [55], we locate the critical temperature and extract the critical exponents for all these order parameters as detailed in Tab. 2. They agree with the existing numerical values very well. We conclude that all occupancies of these states along with the magnetization-like quantity $M_{\mathrm{map}}$ can serve as alternative order parameters for $q=4$ Potts model as well.

### v2

@@ III.1 $q=3$ Potts model on a square lattice
Table 1: Critical properties of the three-state Potts model at $d=2$ on a square lattice, obtained using the alternative order parameters $M_{\mathrm{max}}$, $M_{\mathrm{med}}$, and $M_{\mathrm{min}}$ and quasiorder parameter $M_{\mathrm{map}}$. The critical temperature $T_{c}$, scaling dimension $\Delta_{\sigma}$, and critical exponent $\nu$, along with their standard errors, are estimated via finite-size scaling within the interval $T\in(0.920,1.046)$ with step size $\Delta T=0.002$. Theoretical values from the literature are included for comparison [51].
For the $q=3$ case, the mapping of three spin components takes the form
$\{1,2,3\}\mapsto\{-1,1,0\},$ (6)
where we keep the first two components and ignore the third one by setting it as zero. Building on the established understanding that ANNs trained on Ising models learn to detect magnetization patterns, i.e., the sum of all spin values, we define a corresponding quantity for the mapped Potts model:
$M_{\mathrm{map}}=\frac{|N_{1}-N_{2}|}{N}.$ (7)
$M_{\mathrm{map}}$ represents a natural generalization of order parameter for the Ising model. For $q=2$ Potts model, $M_{\mathrm{map}}$ is equivalent to $M_{\mathrm{max}}$ in Eq. (3) as demonstrated by Eq. (5). Our finite-size scaling analysis (shown in Fig. A2 in the Appendix) demonstrates that $M_{\mathrm{map}}$ successfully captures critical behavior using the known critical exponents $\nu$ and $\Delta_{\sigma}$, confirming its validity as an order parameter. In this work, we are concerned with the static critical properties of Potts model in equilibrium and notably $M_{\mathrm{map}}$ could be viewed as a static version of the alternative order parameter for Potts model at an early stage of the time evolution of dynamics at non-equilibrium [17]. It is also worth noting that $M_{\mathrm{map}}$ has the same form of the order parameter for $Z_{5}$ model, aka five-state clock model [45].
To relate $M_{\mathrm{max}}$ and $M_{\mathrm{map}}$ theoretically, we can make use of the minimal state occupancy in each configuration $\mathrm{min}\{N_{i}\}/N$. As we know, $\mathrm{min}\{N_{i}\}/N$ also tends to $1/q$ when $T>T_{c}$. While at low $T$, this occupancy is suppressed by the dominant state. So, here we construct a quantity $M_{\mathrm{min}}$ by replacing the occupancy of the most state in each configuration $\mathrm{max}\{N_{i}\}/N$ in Eq. (3) with the least one, $\mathrm{min}\{N_{i}\}/N$:
$M_{\mathrm{min}}=\frac{q\frac{\mathrm{min}\{N_{i}\}}{N}-1}{q-1}.$ (8)
We assume that in one configuration, the occupancy of three states descends as $a,b,c$ with $a+b+c=1$. With the Z(3) symmetry, we could have $A_{3}^{3}=6$ cases of how these three occupancies are assigned to these three states. After smearing a specific state of these three ones with the mapping procedure, the average $M_{\mathrm{map}}$ of these six configurations equals 2(ac)/32(a-c)/3. Since $M_{\mathrm{max}}=(3a-1)/2$ and $M_{\mathrm{min}}=(3c-1)/2$, as in Eqs. (3) and (8), the equilibrium averages satisfy the fundamental relationship
$\bar{M}_{\mathrm{map}}=\frac{4}{9}(\bar{M}_{\mathrm{max}}-\bar{M}_{\mathrm{min}}),$ (9)
where the overline signifies the ensemble average. This is also confirmed numerically in the upper panel of Fig. 1, where $M_{\mathrm{max}}$, $M_{\mathrm{min}}$, and $M_{\mathrm{map}}$ as functions of temperature $T$ with lattice size $L=96$ are shown for visual comparison and the curve of $M_{\mathrm{map}}$ labeled with red triangles and that of $\frac{4}{9}(M_{\mathrm{max}}-M_{\mathrm{min}})$ labeled with gray dashed line collapse on top of each other. In the middle and lower panels of Fig. 1, we show the standard error of mean (SEM) and relative standard error (RSE), i.e., the ratio of SEM over mean, of these quantities, respectively. Due to the high statistics in this work (100 000 samples), the standard errors of mean are too small to be visualized along with their mean values in the upper panel. One can observe that below $T_{c}$ the fluctuations of $M_{\mathrm{map}}$ are much larger than those of other quantities, particularly $\frac{4}{9}(M_{\mathrm{max}}-M_{\mathrm{min}})$, which are mainly due to the non-thermal fluctuations between those aforementioned six symmetric configurations for $M_{\mathrm{map}}$ arising from the different treatment on all Potts states.
One can easily understand these non-thermal fluctuations in $M_{\mathrm{map}}$ at $T=0$ limit, where one expects full occupation of one of the Potts states and zero occupation of the other states, i.e., $a=1$ and $b=c=0$. For $q=3$, there are $A_{3}^{3}/A_{2}^{2}=3$ equally likely possibilities,
$(N_{1},N_{2},N_{3})/N=(1,0,0),(0,1,0),(0,0,1).$ (10)
For these three configurations, $M_{\mathrm{map}}$ takes the values 1, 1, and 0, despite that all three configurations are ordered to the same extent. This leads to a mean $\bar{M}_{\mathrm{map}}=2/3$, a standard deviation (SD) $M_{\mathrm{map}}^{\mathrm{SD}}$ = $\sqrt{2}/3$, and a standard error of the mean $M_{\mathrm{map}}^{\mathrm{SEM}}$$M_{\mathrm{map}}^{\mathrm{SD}}/\sqrt{100\,000}=0.00149$. Note that these predictions for $T=0$ are quite compatible with the simulation results in the upper two panels of Fig. 1. For general $q$ states, $\bar{M}_{\mathrm{map}}=2/q$, $M_{\mathrm{map}}^{\mathrm{SD}}$ = $\frac{\sqrt{2(q-2)}}{q}$, and $M_{\mathrm{map}}^{\mathrm{SEM}}$$M_{\mathrm{map}}^{\mathrm{SD}}/\sqrt{n}$, where $n$ is the total sample number.
Note that the fluctuations of $M_{\mathrm{map}}$, quantified by its SEM, do not behave like a susceptibility with a clear peak near $T_{c}$ for finite system size. In this sense, $M_{\mathrm{map}}$ can only be regarded as a quasiorder parameter. However, as $T$ approaches $T_{c}$ from above in the second panel of Fig. 1, the SEM for $M_{\mathrm{map}}$ rises steeply, as if approaching a peak, similar to the behavior of $M_{\mathrm{max}}$ and $M_{\mathrm{min}}$. We interpret this as a susceptibility-like enhancement that, however, is difficult to resolve for $T<T_{c}$ because of additional non-thermal fluctuations inherent to the definition of $M_{\mathrm{map}}$, which increase with decreasing temperature, as the system orders increasingly, and become strongest at $T=0$.
An even simpler quantity than $M_{\mathrm{map}}$ is
$M_{1}=\frac{|q\frac{N_{1}}{N}-1|}{q-1}.$ (11)
It reduces to the Ising order parameter when $q=2$ due to $|\frac{2N_{1}}{N}-1|=\frac{|N_{1}-N_{2}|}{N}$ and is expected to behave as an order parameter in general for $q$-state Potts model. However, it does not treat all Potts states equally and therefore introduces non-thermal fluctuations as $M_{\mathrm{map}}$. This quantity is of interest since it only involves one of the Potts states and provides motivation for ANN studies to extract Potts critical properties from simplified configurations that only retain the occupancy of one of the $q$ states. We include the corresponding results for $M_{1}$ in Fig. 1. Interestingly, the mean values of $M_{1}$ coincide with $M_{\mathrm{map}}$ and its $T=0$ limit can be obtained analytically as $2/q$, consistent with that of $M_{\mathrm{map}}$ as well. The SEM of $M_{1}$, with the limit $\frac{q-2}{q\sqrt{q-1}}\frac{1}{\sqrt{n}}=0.000\,75$ at $T=0$, is smaller than that of $M_{\mathrm{map}}$. More generally, we find that any quantity, such as $M_{\mathrm{map}}$ and $M_{1}$, which treats ordering in a subset of Potts states differently from the others, will have non-thermal fluctuations that grow as $T$ decreases and may obscure the susceptibility peak associated with the thermal fluctuations near the critical point.
$M_{\mathrm{med}}=\frac{q\frac{\mathrm{medium}\{N_{i}\}}{N}-1}{q-1}.$ (12)
Due to the constraint that $M_{\mathrm{max}}+M_{\mathrm{med}}+M_{\mathrm{min}}=1$, only two of them are independent. We therefore conjecture that $M_{\mathrm{med}}$ could serve as an alternative order parameter as well. The temperature dependence of $|M_{\mathrm{min}}|$ and $|M_{\mathrm{med}}|$ and their associated finite-size scaled results are shown, respectively, in Figs. A3 and A4 in the Appendix with the known values of $\nu$ and $\Delta_{\sigma}$, which validates that $|M_{\mathrm{min}}|$ and $|M_{\mathrm{med}}|$ could also serve as alternative order parameters. There we take the absolute value of $M_{\mathrm{min}}$ and $M_{\mathrm{med}}$ so that they increase monotonically as temperature decreases and reach a maximum at the ground state with $T=0$, similar to the traditional order parameter $M_{\mathrm{max}}$. Any order parameter that is symmetric in the occupancies of the $q$ states, such as $M_{\mathrm{max}}$, $M_{\mathrm{min}}$, or $M_{\mathrm{med}}$, takes the same value for each of the three ground-state configurations, and therefore exhibits no fluctuations at $T=0$. Specifically, at this limit one finds $M_{\mathrm{max}}=1$, $M_{\mathrm{min}}=M_{\mathrm{med}}=-1/2$, with both the SD and SEM vanishing. These predictions labeled by the short dashed lines are also quite compatible with the simulation results in the upper two panels of Fig. 1. From the lower panel of Fig. 1, it is obvious that although $M_{\mathrm{max}}$, $M_{\mathrm{med}}$, and $M_{\mathrm{min}}$ all could serve as the order parameters of Potts model, which is consistent with the model’s universality class and symmetry, they are not completely equivalent in the sense of their relative standard error, and $M_{\mathrm{min}}$ exhibits the minimal relative standard error for the whole critical temperature interval of interest.
With the FSS procedure [30], we locate the critical temperature and extract the critical exponents for all these order parameters $M_{\mathrm{max}}$, $M_{\mathrm{med}}$, and $M_{\mathrm{min}}$ and quasiorder parameter $M_{\mathrm{map}}$ as detailed in Table 1. They agree with the theoretical values very well and we conclude that all occupancies of these states along with the magnetization-like quantity $M_{\mathrm{map}}$ can serve as (quasi)order parameters for $q=3$ Potts model. In terms of quantifying order, the suppression of secondary or minimally occupied spin states is directly complementary to the dominance of the majority spin type. This suppression reflects the onset of symmetry breaking and forms the conceptual basis for proposing these quantities as alternative order parameters. While these quantities may be mathematically related to $M_{\mathrm{max}}$, they represent a distinct and meaningful perspective, i.e., focusing on the elimination of minority components rather than the rise of the dominant one. This conceptual shift provides complementary insight into the nature of order in the Potts models and helps to elucidate the features learned by machine-learning models that implicitly capture phase transitions.
It would be interesting to check whether the above observations and conclusion are valid for other $q$-state Potts models. In Ising model, we have two occupancies and only one is independent, which is a trivial case. So, we will turn to $q=4$ Potts model in next subsection.
@@ III.2 $q=4$ Potts model on a square lattice
Figure 2: (Color online) The mean (upper), standard error of mean (SEM) (middle), and relative standard error (RSE) (lower) of the alternative order parameters $M_{\mathrm{max}}$, $M_{\mathrm{max}}^{\mathrm{sec}}$, $M_{\mathrm{min}}^{\mathrm{sec}}$, $M_{\mathrm{min}}$, and quasiorder parameter $M_{\mathrm{map}}$ as functions of temperature $T$ for the four-state Potts model at $d=2$, simulated on a square lattice of size $L=96$. The relation $M_{\mathrm{map}}=\frac{1}{8}(3(M_{\mathrm{max}}-M_{\mathrm{min}})+M_{\mathrm{max}}^{\mathrm{sec}}-M_{\mathrm{min}}^{\mathrm{sec}})$ is confirmed by the overlap of the curves in the upper panel. The blue horizontal dashed lines mark zero to indicate the sign, while the red vertical dashed lines denote the theoretical critical temperature $T_{c}$.
For the $q=4$ case, besides the most and least states, we also have the second most and second least states for each configuration; therefore, we define
@@
and $M_{\mathrm{map}}$ is defined the same as in Eq. (7). As the analysis for three-state Potts model, we assume that in one configuration, the occupancy of four states descends as $a,b,c,d$ with $a+b+c+d=1$. With the Z(4) symmetry, we could have $A_{4}^{4}=24$ cases of how these four occupancies are assigned to these four states. After smearing two specific states of these four ones with the mapping procedure, the average $M_{\mathrm{map}}$ of these 24 configurations equals (3(ad)+(bc))/6(3(a-d)+(b-c))/6. Since $M_{\mathrm{max}}=(4a-1)/3$, $M_{\mathrm{min}}=(4d-1)/3$, $M_{\mathrm{max}}^{\mathrm{sec}}=(4b-1)/3$, and $M_{\mathrm{min}}^{\mathrm{sec}}=(4c-1)/3$, as in Eqs. (3), (8), (13), and (14), the equilibrium averages satisfy the fundamental relationship
@@
where the overline signifies the ensemble average. This is also confirmed numerically in the upper panel of Fig. 2, where $M_{\mathrm{max}}$, $M_{\mathrm{max}}^{\mathrm{sec}}$, $M_{\mathrm{min}}^{\mathrm{sec}}$, $M_{\mathrm{min}}$, $M_{\mathrm{map}}$, and $M_{1}$ as functions of temperature $T$ with lattice size $L=96$ are shown for visual comparison and the curves of $M_{\mathrm{map}}$ labeled with red triangles, $M_{1}$ labeled with yellow dots, and that of $\frac{1}{8}(3(M_{\mathrm{max}}-M_{\mathrm{min}})+M_{\mathrm{max}}^{\mathrm{sec}}-M_{\mathrm{min}}^{\mathrm{sec}})$ labeled with gray dashed line collapse on top of each other. In the middle and lower panels of Fig. 2, we show the SEM and RSE of these quantities, respectively. Similar to $q=3$ case, one can observe that below $T_{c}$ the fluctuations of $M_{\mathrm{map}}$ and $M_{1}$ are much larger than those of other quantities, particularly $\frac{1}{8}(3(M_{\mathrm{max}}-M_{\mathrm{min}})+M_{\mathrm{max}}^{\mathrm{sec}}-M_{\mathrm{min}}^{\mathrm{sec}})$, which are mainly due to the non-thermal fluctuations between those aforementioned 24 symmetric configurations for $M_{\mathrm{map}}$ arising from the different treatment on all Potts states.
We estimate the non-thermal fluctuations in $M_{\mathrm{map}}$ and $M_{1}$ at $T=0$ limit where $a=1$ and $b=c=d=0$. For $q=4$, there are $A_{4}^{4}/A_{3}^{3}=4$ equally likely possibilities,
$(N_{1},N_{2},N_{3},N_{4})/N$
$=(1,0,0,0),(0,1,0,0),(0,0,1,0),(0,0,0,1)$ (17)
For these four configurations, $M_{\mathrm{map}}$ takes the values 1, 1, 0, and 0, despite that all four configurations are ordered to the same extent. This yields $\bar{M}_{\mathrm{map}}=1/2$, $M_{\mathrm{map}}^{\mathrm{SD}}$ = $1/2$, and $M_{\mathrm{map}}^{\mathrm{SEM}}$$M_{\mathrm{map}}^{\mathrm{SD}}/\sqrt{100\,000}=0.001\,58$. While $M_{1}$ takes the values 1, 1/3, 1/3, and 1/3, leading to $\bar{M}_{1}=1/2$, $M_{1}^{\mathrm{SD}}$ = $\sqrt{1/12}$, and $M_{1}^{\mathrm{SEM}}$$M_{1}^{\mathrm{SD}}/\sqrt{100\,000}=0.000\,91$. These $T=0$ predictions are fully consistent with the general $q$-state results discussed in the previous subsection and are quite compatible with the simulation data shown in the upper two panels of Fig. 2. Again, the fluctuations of $M_{\mathrm{map}}$ and $M_{1}$, quantified by their SEM, do not behave like a susceptibility with a peak near the $T_{c}$ for finite system size. In this sense, $M_{\mathrm{map}}$ and $M_{1}$ can only be regarded as quasiorder parameters.
Order parameters that are symmetric in the occupancies of the $q$ states, such as $M_{\mathrm{max}}$, $M_{\mathrm{min}}$, $M_{\mathrm{max}}^{\mathrm{sec}}$, or $M_{\mathrm{min}}^{\mathrm{sec}}$, take the same value for each of the four ground-state configurations, and therefore exhibit no fluctuations at $T=0$. Specifically, at this limit one finds $M_{\mathrm{max}}=1$, $M_{\mathrm{min}}=M_{\mathrm{max}}^{\mathrm{sec}}=M_{\mathrm{min}}^{\mathrm{sec}}=-1/3$, with both the SD and SEM vanishing. These predictions labeled by the short dashed lines are also quite compatible with the simulation results in the upper two panels of Fig. 2.
(Quasi)order parameter $T_{c}$ $\Delta_{\sigma}$ $\nu$
@@
[…]
Theoretical value $\ln{(1+\sqrt{4})}\sim 0.9102 1$/8=0.1251/8=0.125 $2/3~\sim 0.667$
Table 2: Critical properties of the four-state Potts model at $d=2$ on a square lattice, obtained using the alternative order parameters $M_{\mathrm{max}}$, $M_{\mathrm{max}}^{\mathrm{sec}}$, $M_{\mathrm{min}}^{\mathrm{sec}}$, and $M_{\mathrm{min}}$ and quasiorder parameter $M_{\mathrm{map}}$. The critical temperature $T_{c}$, scaling dimension $\Delta_{\sigma}$, and critical exponent $\nu$, along with their standard errors, are estimated via finite-size scaling within the interval $T\in(0.840,0.966)$ with step size $\Delta T=0.002$. For comparison, we include Monte-Carlo numerical estimates of $\Delta_{\sigma}$ and $\nu$ that neglect higher-order corrections and mark them with * [34], along with the theoretical values from the literature [51].
The temperature dependence of $M_{\mathrm{max}}$, $M_{\mathrm{max}}^{\mathrm{sec}}$, $|M_{\mathrm{min}}^{\mathrm{sec}}|$, $|M_{\mathrm{min}}|$, and $M_{\mathrm{map}}$ and their associated finite-size scaled results are shown, respectively, in Figs. A5-A9 in the Appendix with the known Monte-Carlo numerical estimates of $\nu=0.722$ and $\Delta_{\sigma}=0.128$ [34], which validates that all of these quantities could serve as alternative (quasi)order parameters. There we take the absolute value or reverse the sign of $M_{\mathrm{min}}$, $M^{\mathrm{sec}}_{\mathrm{min}}$ and $M^{\mathrm{sec}}_{\mathrm{max}}$ so that they increase monotonically as temperature decreases and reach a maximum at the ground state with $T=0$, similar to the traditional order parameter $M_{\mathrm{max}}$. From the lower panel of Fig. 2, it is obvious that although $M_{\mathrm{max}}$, $M_{\mathrm{max}}^{\mathrm{sec}}$, $M_{\mathrm{min}}^{\mathrm{sec}}$, and $M_{\mathrm{min}}$ all could serve as the order parameters of Potts model, they are not completely equivalent in the sense of their relative standard error, and same as $q=3$ case, $M_{\mathrm{min}}$ exhibits the minimal relative standard error for the whole critical temperature interval of interest. With the FSS procedure [30], we locate the critical temperature and extract the critical exponents for all these (quasi)order parameters as detailed in Table 2. They agree with the existing numerical estimates of $\nu=0.722$ and $\Delta_{\sigma}=0.128$ [34] very well. The deviation from the theoretical values of $\nu=2/3$ and $\Delta_{\sigma}=1/8$ [51] arises from the omission of the higher-order corrections. We conclude that all occupancies of these states along with the magnetization-like quantity $M_{\mathrm{map}}$ can serve as alternative (quasi)order parameters for $q=4$ Potts model as well.